OpenStudy (anonymous):

1.)The perimeter of a rectangle is 52.It's length is 2 more than twice the measure of it's width. Find it's dimension. 2.)If the diameter of a circle is the same as the radius of another circle, what is the ratio of the circumference of the larger circle to the circumference of the smaller ?

OpenStudy (anonymous):

L = 2 + 2W So, 2L + 2W = 52 4 + 4W + 2W = 52 6W = 48 Find W from here and then Find L..

OpenStudy (anonymous):

I don't get it

OpenStudy (anonymous):

What you did not get tell me??

OpenStudy (anonymous):

ah .. i know ... 6W=48 divided by 6 so the width is 8 and then the L .. how to find it ?

OpenStudy (diyadiya):

L = 2 + 2W We know the value of W, so substitute it and find L

OpenStudy (anonymous):

Here, just put the values W = 8 (you found) L = 2 + 2*8 L = 2 + 16 Add it..

OpenStudy (anonymous):

i have a question: why is it that it is 4W eh 2(2+2W) =4+4W right ?

OpenStudy (diyadiya):

L = 2 + 2W 2L + 2W = 52 2(2+2W) + 2W = 52 4 + 4W + 2W = 52 6W = 48

OpenStudy (anonymous):

ah ... i thought it was in width !! SORRY :{

OpenStudy (diyadiya):

he just substituted L with (2+2W)

OpenStudy (anonymous):

See, L = 2 + 2*W Other equation is: \(2\color{green}L + 2W = 52\) \(2\color{green}{(2 + 2W)} + 2W = 52\) This is the way I solved it..

OpenStudy (anonymous):

then how to find the dimension?

OpenStudy (diyadiya):

W is Width & L is Length You already found out W (that is Width) Now plug in the value of W in the equation L = 2 + 2W

OpenStudy (anonymous):

the answer is 16

OpenStudy (anonymous):

I have actually solved for L too.. Can you check there??

OpenStudy (diyadiya):

Oh, Yes! Sorry, i missed that post @waterineyes

OpenStudy (anonymous):

Check my post above I have solved for L too.. Can you tell me the value of L now??

OpenStudy (anonymous):

16

OpenStudy (anonymous):

ohh ... its 18 pala

OpenStudy (anonymous):

Yes you are right..

OpenStudy (anonymous):

then what is the dimension ?

OpenStudy (anonymous):

See, when question asks for Dimensions then it is asking for values of Length and Width.. You have found both so that will be your answers.. You can write it as: Dimensions are: Length = Width =

OpenStudy (anonymous):

ah .. so how about in number 2 ?

OpenStudy (anonymous):

Can you tell me what is the formula for Circumference of a Circle??

OpenStudy (anonymous):

2r ?

OpenStudy (anonymous):

NO..

OpenStudy (anonymous):

2(pi)r

OpenStudy (anonymous):

\(2r\) is Diameter \(2 \pi r\) is Circumference..

OpenStudy (anonymous):

Let the other circle has diameter = r Let the diameter of first circle = 2x So, 2x = r So, \[\frac{2 \pi r}{\pi x} = \frac{2r}{x} = \frac{4r}{2x} = \frac{4r}{r} = 4\]

OpenStudy (anonymous):

where did you get the 4r/2x ??

OpenStudy (anonymous):

I have just multiply numerator and divide by 2..

OpenStudy (anonymous):

ah

OpenStudy (anonymous):

the answer is 4 ?

OpenStudy (anonymous):

The question is asking for ratio.. So, 4:1

OpenStudy (anonymous):

ah :)) thank you so much .... xDD

OpenStudy (anonymous):

Welcome dear..

OpenStudy (anonymous):

ahm ... question : where does the x go ? in 4r/2x ?

OpenStudy (anonymous):

@waterineyes

OpenStudy (anonymous):

See, what is the value of 2x?? 2x = r I have written above.. Got??

OpenStudy (anonymous):

no , in number 1

OpenStudy (anonymous):

In number 1 ??/ where are you pointing??

OpenStudy (anonymous):

ayyy ... i get it !!

OpenStudy (anonymous):

Sure???

OpenStudy (anonymous):

yep yep :)

OpenStudy (anonymous):

So we are done?? Right??

OpenStudy (anonymous):

yep yep :)

OpenStudy (anonymous):

Do solve problems like these two you will surely get command over them... Take care..